OpenStudy (baby456):

Is the function f(x)=1/x even, odd,or neither even or odd? justify your answer?

zepdrix (zepdrix):

\[\large\rm \color{orangered}{f(x)=\frac{1}{x}}\]Can you tell what happens when we plug a negative x into the function?\[\large\rm f(-x)=?\]

OpenStudy (baby456):

we get -1

zepdrix (zepdrix):

Mmmmm no. We didn't plug -1 into the function, we plugged -x in place of x.

OpenStudy (baby456):

i dont know i gues -x

zepdrix (zepdrix):

\[\large\rm f(-x)=\frac{1}{(-x)}\]Replacing x with -x gives us this, yes?

zepdrix (zepdrix):

The brackets are not necessary I suppose

OpenStudy (baby456):

yes

zepdrix (zepdrix):

\[\large\rm f(-x)=-\color{orangered}{\frac{1}{x}}\]Let's pull the negative out front, and recognize that this 1/x is the f(x) function that we started with,\[\large\rm f(-x)=-\color{orangered}{f(x)}\]

zepdrix (zepdrix):

Odd functions have this property, f(-x) = -f(x). So we've shown that the function is `odd`.